Task Processing Time (hrs. 13Use CPM to solve the project management network below using the processing times contained in the accompanying table. Then ma
Answer #1

Answer:

step 1 find early start, and early finish from the start of the project for each activity
step 2 Find Maximum EF and keep it as the LF of last activities. Max of EF's is 56
step 3 calculate LS for each activity after calculating its EF
Activity Duration, D Early start, ES=Max of early finish of preceding activities Early finish, EF = ES + D Late finish, LF= Min of LS of successor activities Late start, LS= LF - D Total slack= LF-EF Critical activity, activities with 0 slack time
A 13 0 13 13 0 0 Yes
B 3 13 16 26 23 10 No
C 13 13 26 26 13 0 Yes
D 19 26 45 45 26 0 Yes
E 11 45 56 56 45 0 Yes
F 8 45 53 56 48 3 No

1. LF for task D= 45

2. Project completion time= 56

3.Critical path= A-C-D-E

4. Slack of task B= 10

Please ask, if you have any doubts through the comment section. Do rate the answer